OpenStudy (anonymous):

Solve each system using elimination. 3x+3y=27 x-3y=-11 9. 2x+4y=22 2x-2y=-8 11. 5x-y=0 3x+y=24 Please show your work!! Thank you!!

OpenStudy (usukidoll):

first one add the y's and solve for x.. then plug back the x to find the y

OpenStudy (anonymous):

3x + 3y = 27 x - 3y = -11 3y cancels out 3x + x = 27 + -11 4x = 16 there x = 4. Sub that into equation (2).... x -3y = -11 4- 3y = -11 -3y = -15 there fore y = 5. x =4, y =5

OpenStudy (anonymous):

Thank you soo much @BizPro Can you help me with the others??

OpenStudy (anonymous):

@ganeshie8 @satellite73 @jim_thompson5910 @Zarkon Please help!!

OpenStudy (anonymous):

I need help with the last two problems, 9 and 11

jimthompson5910 (jim_thompson5910):

for #9, my hint is to multiply everything in equation 2 by 2 to go from this 2x-2y=-8 to this 4x-4y=-16 ------------------------ we now have this system 2x+4y=22 4x-4y=-16 notice how we haev +4y in the top and -4y in the bottom. They add to 0y, aka 0, and go away and the y terms are eliminated when you add the equations of this new system 2x+4y=22 4x-4y=-16 ---------- 6x + 0y = 6 So 6x = 6 ----> x = ????

OpenStudy (anonymous):

x=1?

jimthompson5910 (jim_thompson5910):

Good, now take that value of x and plug it into any equation that has x and y in it. From there solve for y

jimthompson5910 (jim_thompson5910):

For instance, 2x+4y=22 2*1+4y=22 ... plug in x = 1 2 + 4y = 22 ... ... ... y = ???

OpenStudy (anonymous):

Does y=-4??

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

I am confused..

jimthompson5910 (jim_thompson5910):

2 + 4y = 22 4y = 22 - 2 4y = 20 y = ???

OpenStudy (anonymous):

In the first equation..2x+4y=22 we got that x=1. Not to figure out why, I must plug in x into 2x-2y=-8, right? I plugged it into that equation and my answer came out to be y=-4

jimthompson5910 (jim_thompson5910):

2x-2y=-8 2(1)-2y=-8 2 - 2y = -8 -2y = -8 - 2 -2y = -10 y = ???

jimthompson5910 (jim_thompson5910):

No matter which equation you pick, you should get the same answer for y.

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

closer, but no

jimthompson5910 (jim_thompson5910):

-2y = -10 y = -10/(-2) y = 5 ... it's positive since the negatives cancel

jimthompson5910 (jim_thompson5910):

4y = 20 y = 20/4 y = 5 ... same answer

OpenStudy (anonymous):

Okay, can you please help me with the next question?

jimthompson5910 (jim_thompson5910):

Add the equations to get 5x-y=0 ... notice how the y terms are equal yet opposite 3x+y=24 ... notice how the y terms are equal yet opposite --------- 8x + 0y = 24 8x = 24 ----> x = ???

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

Now use that to find y

OpenStudy (anonymous):

Am I doing this right? 3(3)+y=24 9+y=24 9y=24

jimthompson5910 (jim_thompson5910):

9+y does NOT turn into 9y

jimthompson5910 (jim_thompson5910):

9y means "9 times y" and it's completely different from 9+y

OpenStudy (anonymous):

Okay...So then how would I do it? I am lost right now.

jimthompson5910 (jim_thompson5910):

hint: 9+y is the same as y + 9

jimthompson5910 (jim_thompson5910):

so 9+y=24 is the same as y + 9 = 24

OpenStudy (anonymous):

oh..do I switch around the equation? y=24+9?

jimthompson5910 (jim_thompson5910):

no, you need to move that +9 over

jimthompson5910 (jim_thompson5910):

so what must you do to both sides?

OpenStudy (anonymous):

divide or subject?

jimthompson5910 (jim_thompson5910):

which operation undoes addition?

OpenStudy (anonymous):

subtract, sorry. I mixed up the word with subject

jimthompson5910 (jim_thompson5910):

it's ok

jimthompson5910 (jim_thompson5910):

which operation undoes addition?

OpenStudy (anonymous):

now Y equals 24-9

OpenStudy (anonymous):

y=15

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

wow! Thank you soo much for your help! Can you help me with a few more?

jimthompson5910 (jim_thompson5910):

I can help with one more

OpenStudy (anonymous):

Okay. 12x+2y=90 6x+2y=60

jimthompson5910 (jim_thompson5910):

notice if you add the equations, the y terms will be 2y+2y = 4y, which is NOT 0y or 0 like we want. We can fix this by changing the signs of each term in either equation. So change the signs of each term in the second equation to get -6x - 2y = -60. Now add the equations 12x + 2y = 90 -6x - 2y = -60 -------------- 6x + 0y = 30 Notice how the y terms are gone now leaving us with 6x = 30 x = ??

OpenStudy (anonymous):

x=5

jimthompson5910 (jim_thompson5910):

Now use it to find y

OpenStudy (anonymous):

y=15

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

I have 3 more of the equations left, can I work them through and you just watch to make sure I am doing them right?

jimthompson5910 (jim_thompson5910):

Sure I can do that. Go ahead

OpenStudy (anonymous):

Thanks. Okay, the first one is 2x+3y=9 x+5y=8

OpenStudy (anonymous):

2x+3y=9 x +5y=8 _________ 2x+8y=17 Is that right so far?

jimthompson5910 (jim_thompson5910):

well first off, 2x + x is 3x (not 2x) and notice how 3y + 5y = 8y which is not 0y or 0 We want either the x or y terms to be eliminated. This doesn't happen here.

jimthompson5910 (jim_thompson5910):

What can you do to fix this?

OpenStudy (anonymous):

Okay, let me try again. 2x+3y=9 x -5y=8 _________ 3x+0y=17 Now is that right?

jimthompson5910 (jim_thompson5910):

now 3y + (-5y) = -2y which is still not 0y

jimthompson5910 (jim_thompson5910):

also, you need to do the same thing for all terms in the second equation (you cannot just change the sign of one term)

OpenStudy (anonymous):

I don't know what to do now.

jimthompson5910 (jim_thompson5910):

we have 2x in the top equation and an 'x' in the bottom equation why not multiply that x in the bottom equation by -2 to get -2x now notice how 2x + (-2x) = 2x - 2x = 0x = 0

jimthompson5910 (jim_thompson5910):

So what you have to do is multiply everything in equation 2 by -2 to get x+5y=8 -2*x + (-2)(5y) = -2*8 -2x - 10y = -16

jimthompson5910 (jim_thompson5910):

This means the system 2x + 3y = 9 x +5y = 8 turns into 2x + 3y = 9 -2x - 10y = -16

OpenStudy (anonymous):

And now do I subtract?

jimthompson5910 (jim_thompson5910):

add the equations actually

jimthompson5910 (jim_thompson5910):

since 2x + (-2x) = 0x

OpenStudy (anonymous):

so it would be -4x-13y=-25?

jimthompson5910 (jim_thompson5910):

you're adding the equations incorrectly (see my last post)

OpenStudy (anonymous):

x+13y=-25?

jimthompson5910 (jim_thompson5910):

2 + (-2) is the same as 2 - 2

OpenStudy (anonymous):

I am lost again.

jimthompson5910 (jim_thompson5910):

2x + (-2x) is the same as 2x - 2x

OpenStudy (anonymous):

I understand that, so then why is 0x-13y=-25 wrong?

jimthompson5910 (jim_thompson5910):

3y + (-10y) is not -13y

jimthompson5910 (jim_thompson5910):

9 + (-16) is not -25

OpenStudy (anonymous):

0x=-7y=-7?

jimthompson5910 (jim_thompson5910):

closer, but it should be 0x + (-7y) = -7 which is I'm sure what you meant to say

jimthompson5910 (jim_thompson5910):

so... 0x - 7y = -7 -7y = -7 y = ???

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

Now x=3

jimthompson5910 (jim_thompson5910):

perfect on both

OpenStudy (anonymous):

Okay, can I try another one?

OpenStudy (anonymous):

6x+4y=42 -3x+3y=-6 6x+4y=42 -3x+3y=-6 _____________ 3x+7y=26 Does that look right?

jimthompson5910 (jim_thompson5910):

were any variables eliminated?

OpenStudy (anonymous):

No..so do I subtract instead?

jimthompson5910 (jim_thompson5910):

hint: look at the x terms and try to eliminate them

OpenStudy (anonymous):

is x=5

jimthompson5910 (jim_thompson5910):

Do you see how to get x = 5?

OpenStudy (anonymous):

well it had -9x=-48..I divided it.

jimthompson5910 (jim_thompson5910):

-9x = -48 doesn't lead you to x = 5

OpenStudy (anonymous):

Is it wrong?

jimthompson5910 (jim_thompson5910):

add 6x and -3x to get 3x That is NOT 0x how can we change that -3x into something that adds to 6x to get 0x?

OpenStudy (anonymous):

multiply it

jimthompson5910 (jim_thompson5910):

by what

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so it would be -6x+6y=-12

OpenStudy (anonymous):

y=3